OpenStudy (anonymous):

choose the correct simplification of (4x^3-3x-7)-(3x^3+5x+3)

OpenStudy (anonymous):

a.) 7x^3-2x-4 b.)x^3-8x-10

jimthompson5910 (jim_thompson5910):

First distribute to get (4x^3-3x-7)-(3x^3+5x+3) 4x^3-3x-7-3x^3-5x-3 now combine like terms

OpenStudy (anonymous):

c.) 7x^3+2x-4 d.)x^3+8x+10

OpenStudy (anonymous):

Combine like terms. Let's change the - to a + first. (4x^3-3x-7)+(-3x^3-5x-3) Now, this is simply addition. Let's group the like terms. (4x^3-3x^3)+(-3x-5x)+(-7-3) And combine. x^3-8x-10. Choice B.

OpenStudy (anonymous):

(4x^3-3x-7)-(3x^3+5x+3) should actually be (4x^3-3x-7) + (3x^3+5x+3)

jimthompson5910 (jim_thompson5910):

(4x^3-3x-7) + (3x^3+5x+3) 4x^3-3x-7 + 3x^3+5x+3 now combine like terms

OpenStudy (anonymous):

I got 9x^6+10 I know its wrong though :/

jimthompson5910 (jim_thompson5910):

4x^3 and 3x^3 are like terms this is because they both have the same variable term x^3

jimthompson5910 (jim_thompson5910):

to combine like terms, you just add the numbers in front of the variable terms these are the coefficients, which are 4 and 3 4+3 = 7

jimthompson5910 (jim_thompson5910):

so 4x^3 + 3x^3 = 7x^3

jimthompson5910 (jim_thompson5910):

do the same with other like terms

OpenStudy (anonymous):

okay I think the answer is a

jimthompson5910 (jim_thompson5910):

which is what?

OpenStudy (anonymous):

7x^3-2x-4

jimthompson5910 (jim_thompson5910):

-3x + 5x = ???

OpenStudy (anonymous):

oooo wait its actual 7x^3+2-4

jimthompson5910 (jim_thompson5910):

better, it's 7x^3 + 2x - 4

OpenStudy (anonymous):

thanks :) can you help me with another?

jimthompson5910 (jim_thompson5910):

sure I can do one more

OpenStudy (anonymous):

two groups of campers, group a and group b hiked two different routes from their campsite to a waterfall. group a hiked through vista point to reach the waterfall. group b hiked directly to the waterfall the figure shows the distance they hiked: which simplified expression can be used to find the difference in miles hiked by the two groups? a.)2x-1 b.)2x+1 c.)2x+3 d.)2x+7

jimthompson5910 (jim_thompson5910):

how far did group a hike?

OpenStudy (anonymous):

Total distance for group A would be (x-1) + (2x+5) or 3x + 4. Total distance for group B would just be x + 3. I think the answer is b?

jimthompson5910 (jim_thompson5910):

group A traveled more, so you'll find the difference like so (group A) - (group B) (3x + 4) - (x + 3) 3x + 4 - x - 3 2x + 1 so you are correct

OpenStudy (anonymous):

yay!!! :) thanks I really appreciate your help

jimthompson5910 (jim_thompson5910):

you're welcome